OpenStudy (wwe123):

cosø+sinø-sin2ø=1/2 ,0<=ø>=180 degree

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

sorry it's hard

OpenStudy (anonymous):

sin2 theta = 2sintheta costheta

OpenStudy (anonymous):

factorize

OpenStudy (anonymous):

cos x + sinx - 2 sinx cosx = 1/2 sinx ( 1 - 2 cosx) = (1-2cosx)/2 sinx =1/2 x = 30 degree

OpenStudy (anonymous):

there are more answers

OpenStudy (anonymous):

\[\cos x + \sin x - 2 \sin x \cos x = \frac{1}{2}\]\[\cos x -\frac{1}{2}+ \sin x - 2 \sin x \cos x =0\]\[\frac{1}{2}(2\cos x -1)- \sin x (2 \cos x-1) =0\]\[(2\cos x-1)(\sin x-\frac{1}{2})=0\]according to the zero product property \(\cos x =\frac{1}{2}\) it gives \(x=\frac{\pi}{3}\) or \(\sin x =\frac{1}{2}\) it gives \(x=\frac{\pi}{6},\frac{5\pi}{6}\)